a gun is moving  at a speed of 30m/s fires a bullet at an angle of 30o with a veocity of 150m/s relative to the gun.Find the distance between the gun and the bullet when it hits the ground.(g=10m/s2.)
 

Grade:Select Grade

1 Answers

Arun
5 years ago
 

Vertical component of velocity = 150 sin 30o  = 75 m/sec

Horizontal component of velocity relative to gun = 150 cos 30o = 75√3 m/sec

Horizontal component of velocity relative to ground = 75√3 + 30 ≈ 160m/sec

Time of flight   = (2\times75)/g = 15 s

Range of projectile = 160 × 15 = 2400 m

Distance moved by the gun and projectile = 2400 - 450 = 1950 m.